Please help me solve this equation: 3x-6y=6; x=2
Thank you

Answer by checkley77(12844)

3x-6y=6
-6y=-3x+6
y=-3x/-6+6/-6
y=x/2-1
x=2 then:
y=2/2-1
y=1-1
y=0